ind the base of triangle if area is 30 cm2 and height exceeds the base by 7cm
Answer by checkley77(12844)

A=BH/2
30=B(B+7)/2
30=(B^2+7B)/2
B^2+7B=2*30
B^2+7B=60
B^2+7B-60=0
(B+12)(B-5)=0
B-5=0
B=5 ANS FOR THE BASE.
H=5+7=12 ANS FOR THE HEIGHT.
PROOF:
30=5*12/2
30=60/2
30=30